OpenStudy (anonymous):

which two numbers are next in the sequence: 1,3,7,15?

OpenStudy (anonymous):

you can see, 3-1=2 7-3=2^2 15-7=2^3

OpenStudy (anonymous):

so the differences are in gp, first term is 2, common ratio is 2 .. now try to find out the next numbers

OpenStudy (anonymous):

\[a_0 =1\\ a_n = a_{n-1} + 2^n\\ a_5=?? \]